1. Calculate the freezing point of a 0.2 m CaCl2 aqueous solution, assuming it dissociates completely to form ions in solution
Answer #1

i for CaCl2 has it dissociates into 3 particles. [1 Ca2+ and 2 Cl-]


lets now calculate ΔTf
ΔTf = i*Kf*m
= 3.0*1.86*0.2
= 1.116 oC

This is decrease in freezing point
freezing point of pure liquid = 0.0 oC
So, new freezing point = 0 - 1.116
= -1.116 oC
Answer: -1.12 oC